### v1

# Benchmarking Deep Neural Networks for Modern Recommendation Systems

Abstract
This paper examines the deployment of seven different neural network architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—on three distinct datasets: Retail E-commerce, Amazon Products, and Netflix Prize. It evaluates their effectiveness through metrics such as accuracy, recall, F1-score, and diversity in recommendations. The results demonstrate that GNNs are particularly adept at managing complex item relationships in e-commerce environments, whereas RNNs are effective in capturing the temporal dynamics that are essential for platforms such as Netflix.. Siamese Networks are emphasized for their contribution to the diversification of recommendations, particularly in retail settings. Despite their benefits, issues like computational demands, reliance on extensive data, and the challenge of balancing accurate and diverse recommendations are addressed. The study seeks to inform the advancement of recommendation systems by suggesting hybrid methods that merge the strengths of various models to better satisfy user preferences and accommodate the evolving demands of contemporary digital platforms.
I Introduction
Technological advancements and changes in consumer behavior have driven an unprecedented surge in the digital marketplace in recent years. With an extraordinary 8% increase from the previous year, the first quarter of 2023 alone recorded over 540 million transactions, generating revenue exceeding 41 billion euros [1, 2, 3]. This surge not only emphasizes the dynamic character of e-commerce but also poses a critical question: What are the factors propelling this exponential increase in online transactions, and how can we capitalize on them?
Sophisticated recommendation systems [4, 5, 6] are the foundation of this expansion, as they are essential for improving user engagement and sales conversion rates on online platforms. These systems utilize advanced algorithms to analyze extensive quantities of user data, including preferences, purchase history, and interactions, in order to provide highly personalized product recommendations. This optimizes the purchasing experience by aligning product offerings with individual preferences and significantly influencing overall market trends.
The capacity to provide a wide range of products is a critical characteristic of successful recommendation systems [7, 8, 9, 10]. The user experience is enhanced by the diversity of recommendations, which reduces the risk of informational lock-in, a phenomenon that occurs when a limited focus restricts user exposure to a wider variety of products [11, 12]. By encouraging users to explore new products and broadening the recommended items, these systems potentially increase customer loyalty and satisfaction [13, 14, 15].
However, striking an equilibrium between precision and recommendation diversity remains a formidable challenge [16, 17, 18]. Many systems excel at providing precise recommendations; however, they frequently fail to ensure that these recommendations are sufficiently varied to enhance user engagement and discovery. Additionally, the efficacy of these systems varies substantially across different neural network architectures, each offering unique advantages and constraints in the processing and analysis of user data.
The objective of this investigation is to perform a thorough comparative analysis of seven distinct neural network architectures for item–item recommendation systems and to assess their performance in terms of precision and diversity of recommendations. The architectures examined include Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Autoencoders, Transformers, Neural Collaborative Filtering (NCF), and Siamese Networks.
Through a comprehensive evaluation of these models using three separate datasets, this study aims to identify the architectures that most effectively enhance both the accuracy and the diversity of recommendations in e-commerce. The ultimate goal is to provide a practical and exhaustive guide for selecting and implementing the most effective neural network methodology, thereby contributing to the advancement of the digital marketplace.
The paper is systematically structured as follows: Section 2 provides a thorough assessment of the literature, emphasizing key research and advancements in neural networks for recommendation systems. Section 3 offers a detailed exposition of the methodology, encompassing the experimental design, analytical techniques, and implementation procedures. Section 4 presents the results and discussion, interpreting the findings from the comparative study. Section 5 concludes the study by summarizing the principal insights and examining the broader implications of the research findings.
II Related Works
II-A Overview of Recommendation Systems
Recommendation systems play a crucial role in enhancing user experience on e-commerce platforms, and they manifest in various types [18, 19, 20, 21]. Among the most recognized are collaborative filtering, user-based, and content-based methods. These approaches have been extensively applied in both research and practical settings, particularly within the e-commerce industry [22, 23, 24].
In a 2024 study, Enqi Yu et al. [25] introduced a federated recommendation algorithm called ClusterFedMet, which combines user clustering and meta-learning to enhance efficiency and personalization. Their approach addresses challenges related to Non-IID data in federated learning. Although their method reduces communication overhead and improves personalization, issues related to user privacy and data security remain unresolved.
Similarly, Huiying Shi et al. [26] developed a personalized image aesthetics assessment method integrating Graph Neural Networks (GNNs) and collaborative filtering to model both demographic and visual aesthetic interactions. While highly effective, the model’s adaptability across datasets with differing attribute interactions remains a challenge.
Mohsen Jozani et al. [27] conducted a comprehensive empirical analysis of content-based filtering (CBF) systems in mobile app markets, demonstrating how CBF promotes niche products and supports long-tail markets. The study highlights the positive role of CBF in spreading consumer demand across less popular items.
X.J. Li et al. [28] proposed a hybrid recommendation algorithm integrating sentiment analysis (via LSTMs) with matrix decomposition to enhance accuracy. Despite improved performance, computational efficiency and scalability issues remain.
II-B Neural Network Architectures in Recommender Systems
Neural networks have become a cornerstone of modern AI due to their ability to model nonlinear and complex data relationships [29, 30, 31]. Their integration into recommendation systems has transformed the field by enabling the processing of vast amounts of user interaction data [32].
In 2023, Bin Deng et al. [33] introduced a high-performance auditory perception architecture using spiking neural networks (SNNs) implemented on FPGAs. Despite achieving strong performance, issues related to complexity and power consumption limit scalability.
Jiaqi Yan et al. [34] applied neural architecture search (NAS) to design efficient SNNs, introducing a branchless spiking supernet that reduces computational overhead. However, further optimization of the search algorithm is still needed.
Ebrahim Parcham et al. [35] proposed HybridBranchNet, a CNN architecture optimized across depth, width, and resolution. The architecture offers enhanced performance but requires careful parameter tuning.
II-C Accuracy in Recommendation Systems
A 2024 study by Lei Hou and Yichen Huang [36] analyzed the impact of recommendation list length on accuracy and diversity using datasets such as Steam, MovieLens, and Amazon. The work highlights a trade-off between longer lists (more diversity) and accuracy (reduced retrieval precision).
Muhammad Umar et al. [37] studied financial restatements and their impact on sell-side recommendation accuracy. Results show decreased accuracy for buy-and-hold strategies but improved precision for sell recommendations.
II-D Diversity in Recommendation Systems
Diversity ensures that recommendation systems present a wide range of items [38], reducing overspecialization [39] and mitigating echo chamber effects [40]. It encourages users to explore unfamiliar items [41, 42].
Dunlu Peng and Yi Zhou [43] proposed LAP-SR, a post-processing framework enhancing long-tail exposure in session-based recommendations using personalized diversity.
Zihao Li et al. [44] introduced Teddy, a sequential recommendation model disentangling interest trends and diversity using TCNs and MLPs. While effective, dual-pathway complexity may hinder real-time deployment.
Huaizhen Kou et al. [45] developed DI-RAR, a diversity-driven approach for API recommendations based on mashup graphs. Computational overhead remains a challenge.
Sofia Morgado Pereira et al. [46] analyzed diversity in earthquake preparedness recommendations across Europe, revealing significant communication inconsistencies.
Alvise De Biasio et al. [47] proposed techniques to optimize recommendations for sensitive users, reducing harmful content exposure while increasing content diversity.
II-E Positioning Our Work in the State of the Art
Recent studies highlight the evolution of neural-network-based recommender systems, including GNN applications [26] and neural architecture optimization [34]. However, these works rarely focus simultaneously on precision and diversity in item-item e-commerce recommendations.
Our work fills this research gap by comparing seven neural network architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—across multiple datasets to evaluate their performance regarding both accuracy and recommendation diversity. Table I present the summary of the state of the art.
TABLE I: Summary of State-of-the-Art Works
Ref Year Approach Advantages Limitations Dataset
[25] 2024 ClusterFedMet Enhances personalization Privacy and security unexplored -
[26] 2024 Graph-based aesthetic assessment Enhanced visual recommendation Requires high-quality images; computationally intensive Image datasets
[27] 2023 Content-based filtering Enhances novelty and diversity Risk of overspecialization MovieLens 1M
[28] 2023 Sentiment-based hybrid RS Improves accuracy via sentiment High computational cost BeerAdvocate, Modcloth, Amazon
[33] 2023 SNN architectures Efficient neuromorphic processing Complex implementation, limited scalability -
[34] 2024 NAS for SNNs Optimized SNN design High computational resources -
[35] 2023 HybridBranchNet Scalable CNN architecture Requires careful tuning Image datasets
[36] 2024 RS network connectivity Enhances navigation Trade-off depth vs breadth Steam, MovieLens, Amazon
[37] 2023 Financial RS accuracy study Higher accuracy in financial settings Limited generalizability BRICS stock markets
[43] 2024 LAP-SR Enhances long-tail exposure Algorithmic complexity E-commerce datasets
[44] 2024 Teddy model Models trend + diversity Dual-path complexity E-commerce datasets
[45] 2023 DI-RAR Captures implicit API requirements High computational overhead API datasets
[46] 2024 Earthquake recommendation diversity Improves public preparedness Limited to seismic domain EU datasets
[47] 2023 Sensitive-user RS Balances influential items Scalability challenges Social networks
Our Approach 2025 Comparison of 7 NN architectures Precision + diversity evaluation Computational cost trade-offs Retail Rocket, Amazon, Netflix
III Method
In this section, we describe the methodological framework employed to evaluate and compare the performance of seven neural network architectures in item–item recommendation systems for e-commerce. Our study relies on three heterogeneous datasets—Amazon, Netflix Prize, and Retail Rocket E-commerce—each offering distinct behavioral and interaction patterns. Raw data from these sources undergoes rigorous preprocessing to ensure consistency, completeness, and suitability for machine learning workflows.
Each of the seven neural architectures—Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Autoencoders, Transformers, Neural Collaborative Filtering (NCF), and Siamese Networks—is then applied to the datasets. The evaluation focuses on each model’s ability to balance precision and diversity in recommendations. Performance is assessed using precision, recall, diversity score, and computational efficiency, allowing a thorough comparison of the strengths and limitations of each architecture in realistic e-commerce scenarios.
This comparative analysis aims to provide insights into which models best enhance recommendation quality, scalability, and adaptability. All data handling procedures strictly follow privacy-preserving principles and ethical standards. Figure 1 presents the general workflow of the proposed approach.
Figure 1: General workflow of the proposed approach.
III-A Data Preprocessing
Before model implementation, an extensive preprocessing phase is conducted to prepare the datasets. This step is essential for ensuring data quality and maximizing the effectiveness of the neural architectures. The preprocessing pipeline includes missing-value handling, normalization, categorical encoding, and dimensionality reduction. The data is then divided into training and testing subsets. The overall procedure is summarized in Algorithm 1.
1
2Step 1: Input: Raw_Data (user interactions, item metadata, auxiliary fields).
3 Step 2: Output: Preprocessed_Data ready for model training and evaluation.
4Step 3: Data Cleaning:
53.1: Remove corrupted or incomplete entries.
63.2: Fill missing values using statistical or model-based imputation.
73.3: Remove duplicate records.
8Step 4: Text Normalization (if applicable):
94.1: Convert all text to lowercase.
104.2: Remove punctuation, symbols, and extra whitespace.
114.3: Apply stemming or lemmatization.
12Step 5: Tokenization and Encoding:
135.1: Tokenize text fields.
145.2: Encode categorical variables (Label Encoding, One-Hot, TF-IDF, or embeddings).
155.3: Build word/item embeddings if required.
16Step 6: Numerical Feature Scaling:
176.1: Apply MinMaxScaler, StandardScaler, or normalization to a common range.
18Step 7: Feature Optimization:
197.1: Perform feature selection (filter, wrapper, or embedded methods).
207.2: Optionally apply dimensionality reduction (PCA, SVD, Autoencoders).
21Step 8: Finalization:
228.1: Merge cleaned, encoded, and scaled variables.
8.2: Output Preprocessed_Data.
Algorithm 1 Data Preprocessing Pipeline
For the GNN-based models, datasets are structurally converted into undirected graphs. Each node represents an item and contains a single attribute (type) to preserve privacy. Edges encode relationships such as co-occurrences or interactions. The graph construction process is detailed in Algorithm 2.
1
2Step 1: Input: Preprocessed_Data containing item interactions, co-occurrences, and item attributes.
3 Step 2: Output: Training graph $G$ and testing graph $G^{\prime}$.
4Step 3: Dataset Splitting:
53.1: Split Preprocessed_Data into 70% Train_Data and 30% Test_Data.
63.2: Initialize two empty graphs $G$ (training) and $G^{\prime}$ (testing).
7Step 4: Node Construction:
84.1: For each item in Train_Data:
9   Insert a node with attribute: type $\rightarrow$ category.
104.2: Repeat the same process for Test_Data to build $G^{\prime}$.
11Step 5: Edge Construction:
125.1: For each user session or order in Train_Data:
13   Add edges between items purchased/viewed together (co-occurrence).
14   Assign edge weight $w_{ij}$ proportional to co-occurrence frequency.
155.2: Apply the same procedure to Test_Data for graph $G^{\prime}$.
16Step 6: Final Output:
6.1: Return training graph $G$ and testing graph $G^{\prime}$ for GNN model usage.
Algorithm 2 Graph Construction for GNN-Based Recommendation
III-B Model Selection
The seven neural architectures were selected to provide a comprehensive evaluation of different deep learning paradigms used in item–item recommendation. Each architecture contributes unique strengths:
  • CNNs: Capture hierarchical patterns in interaction matrices through spatial feature extraction.
  • RNNs: Model sequential dependencies, suitable for temporal user behavior.
  • GNNs: Naturally process graph-structured data representing item relationships.
  • Autoencoders: Learn compact latent representations, useful for dimensionality reduction.
  • Transformers: Capture long-range dependencies using self-attention mechanisms.
  • NCF: Learn non-linear user–item interactions by combining MLPs and generalized matrix factorization.
  • Siamese Networks: Measure item similarity by comparing item pairs, enhancing diversity.
This selection ensures a balanced comparison of accuracy, diversity, scalability, and representational capacity.
III-C Model Implementation
Table II summarizes the implementation details for each model, including tools, configurations, and training parameters, enabling reproducibility.
TABLE II: Implementation Overview
Model Type Software / Tools Configuration Parameters
CNNs TensorFlow, Keras 2D convolutions, ReLU, max-pooling, dropout LR: 0.001; Batch: 128; Epochs: 30
RNNs PyTorch LSTM units with 0.5 dropout Seq. length: 10; LR: 0.01; Batch: 64; Epochs: 50
GNNs DGL, PyTorch Item–item graph neural network LR: 0.005; Epochs: 40
Autoencoders TensorFlow, Keras Symmetric dense layers; sigmoid output LR: 0.001; Batch: 256; Epochs: 50
Transformers HuggingFace Transformers BERT-like encoder, positional encoding, multi-head attention LR: 0.0001; Batch: 32; Epochs: 20
NCF TensorFlow, Keras MLP + generalized matrix factorization LR: 0.0005; Batch: 128; Epochs: 30
Siamese Networks TensorFlow, Keras Dual subnetworks for similarity learning LR: 0.0005; Batch: 64; Epochs: 35
IV Experimentation and Results
Research Questions
To guide our experimental protocol and evaluate the effectiveness of the seven neural network architectures, the following research questions (RQs) have been formulated:
  • RQ1: How do different neural network architectures perform in terms of accuracy across heterogeneous recommendation datasets?
  • RQ2: Which models provide the highest recall and F1-score, and how consistent are these performances across datasets?
  • RQ3: How does the top-$k$ accuracy degrade as $k$ increases, and which models retain the highest performance as the recommendation list expands?
  • RQ4: Which neural network architecture provides the greatest intra-list diversity in top-$k$ recommendations?
  • RQ5: How do accuracy and diversity interact within each model, and which architectures offer the best trade-off between relevance and variety?
These research questions are addressed systematically throughout the following subsections.
IV-A Dataset (RQ1)
In our research, we have chosen three distinct datasets from different e-commerce and streaming platforms to evaluate the strength and broad applicability of our methods. These datasets were preprocessed as outlined in earlier sections.
Retail Rocket Dataset: This dataset includes category trees, item properties, and user behavior logs, providing insights into consumer actions such as visits and checkouts.
Amazon Product Dataset: Sourced from Amazon, it contains product metadata and user reviews, offering rich information on item popularity and user-item interactions.
Netflix Prize Dataset: This dataset includes detailed user ratings for movies, making it suitable for evaluating temporal user preferences and recommendation quality in streaming platforms.
IV-B Evaluation Metrics (RQ1, RQ2)
To assess the performance of the neural networks, we adopt widely recognized evaluation metrics:
Accuracy:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
where $TP$ is the number of true positives, $TN$ is the number of true negatives, $FP$ is the number of false positives, and $FN$ is the number of false negatives.
Accuracy@k:
[…]
where $k$ is the number of top recommended items and “relevant items” are those that match the user’s actual preferences.
Recall:
$\text{Recall}=\frac{TP}{TP+FN}$ (3)
where $TP$ is the number of true positives and $FN$ is the number of false negatives.
F1-Score:
$\text{F1-Score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
where $\text{Precision}=\frac{TP}{TP+FP}$ and $\text{Recall}=\frac{TP}{TP+FN}$.
Intra-list Diversity@k (ILD@k):
$\text{ILD@k}=1-\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\text{similarity}(item_{i},item_{j})$ (5)
where $k$ is the number of recommended items considered, and $\text{similarity}(item_{i},item_{j})$ measures how similar two items are (e.g., based on feature embeddings or user ratings).
IV-C Justification of Model Parameters (RQ1)
The neural network configurations were selected after iterative tuning to achieve optimal performance.
CNNs: Use 2D convolutions with ReLU for nonlinear learning, dropout for regularization, and max-pooling to reduce dimensionality.
RNNs: LSTM units were chosen for capturing temporal patterns. A dropout of 0.5 prevents overfitting.
GNNs: Two graph convolution layers enable learning representations based on user-item connections. A learning rate of 0.005 ensures stable convergence.
Autoencoders: Symmetric encoder-decoder architecture captures latent representations, with sigmoid normalization in the output.
Transformers: Multi-head attention enables modeling long-range dependencies; positional encoding preserves sequence order.
NCF: Combines matrix factorization with MLP to model nonlinear user-item interactions.
Siamese Networks: Designed for similarity learning, providing strong diversity in item-to-item recommendations.
IV-D Overview of Model Performance (RQ1, RQ2)
Performance across the three datasets is summarized in Tables 3–5 and visualized in Figures 2–4.
TABLE III: Performance metrics on Retail Rocket E-commerce dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 80 83 80.5
RNN 83 86 83.5
GNN 88 91 88.5
Autoencoder 73 76 73.5
Transformer 86 88 86.5
NCF 82 81 82.5
Siamese Networks 81 83 81.5
TABLE IV: Performance metrics on Amazon dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 78 79 77.5
RNN 82 85 82.5
GNN 90 93 90.5
Autoencoder 70 73 70.5
Transformer 88 90 88
NCF 85 84 85.5
Siamese Networks 84 86 84.5
TABLE V: Performance metrics on Netflix Prize dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 75 78 75.5
RNN 89 92 89.5
GNN 85 87 85.5
Autoencoder 72 75 72.5
Transformer 83 85 83.5
NCF 80 79 80
Siamese Networks 78 80 78.5
Figure 2: Metrics visualization on Retail Rocket E-commerce dataset
Figure 3: Metrics visualization on Amazon dataset
Figure 4: Metrics visualization on Netflix Prize
IV-E Accuracy and Diversity Evaluation on Top-$k$ (RQ3, RQ4)
Tables 6–12 present accuracy and intra-list diversity for top-$k$ recommendations for each model-dataset pair. These results are visualized in Figure 5.
TABLE VI: CNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 86.12 0 84.76 0 82.66 0
2 81.46 21.34 80.30 21.75 78.22 22.37
3 79.15 24.72 78.08 25.18 76.04 25.88
4 77.98 29.49 76.43 30.04 74.64 31.14
5 75.85 37.27 74.36 37.75 72.46 39.11
6 71.12 42.72 70.15 43.42 67.86 45.12
7 64.68 47.45 63.61 48.14 61.83 49.63
8 61.50 55.10 60.46 56.06 59.14 57.46
9 58.36 61.71 57.54 62.88 55.92 64.97
10 56.21 66.37 55.09 67.23 53.40 69.40
TABLE VII: RNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 89.43 0 88.48 0 93.41 0
2 86.65 20.64 85.61 20.92 89.49 21.33
3 80.51 25.62 79.63 25.98 83.90 26.74
4 77.98 29.39 76.89 29.75 81.31 30.71
5 74.45 32.83 73.70 33.19 77.18 34.30
6 71.12 38.27 70.39 38.82 73.91 39.68
7 65.78 45.45 64.39 46.12 68.28 47.27
8 61.90 49.10 61.19 49.69 64.12 51.12
9 58.72 54.85 57.94 55.59 61.08 56.86
10 57.12 60.97 56.47 61.75 59.59 63.40
TABLE VIII: GNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 95.63 0 98.47 0 88.39 0
2 92.76 19.34 96.12 7.37 85.83 32.43
3 91.11 23.62 94.38 12.49 82.59 36.73
4 87.98 28.19 91.93 18.25 78.29 40.19
5 84.45 33.83 86.43 30.92 74.30 47.32
6 79.12 37.27 81.09 38.12 68.34 52.49
7 77.28 39.45 77.32 44.16 61.95 55.91
8 74.90 42.10 74.31 47.59 55.12 62.03
9 69.42 52.85 70.98 56.42 49.29 67.30
10 66.12 56.97 67.04 59.31 44.21 71.23
TABLE IX: Autoencoder performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 80.15 0 76.14 0 76.14 0
2 75.41 20.54 73.15 21.20 73.15 21.88
3 73.23 25.13 71.03 25.98 71.03 26.86
4 72.54 30.05 70.36 31.13 69.09 32.25
5 71.23 36.13 69.09 37.50 62.74 38.93
6 65.91 39.88 62.74 41.48 56.91 43.14
7 59.82 42.29 56.91 44.07 51.34 45.92
8 57.72 56.87 55.99 59.37 49.32 61.98
9 54.16 59.31 52.34 62.04 48.32 64.89
10 52.67 62.40 51.09 65.40 46.51 68.54
TABLE X: Transformer performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 90.43 0 97.28 0 85.39 0
2 88.63 19.21 94.12 9.97 83.83 28.43
3 83.51 24.67 93.36 11.59 79.59 31.73
4 79.98 28.49 90.93 17.25 75.29 35.19
5 77.32 31.43 87.13 28.42 71.30 39.32
6 73.12 36.27 83.09 35.13 68.34 46.49
7 67.58 44.47 80.42 41.16 63.95 49.91
8 63.90 47.10 77.33 45.55 58.12 55.03
9 60.82 52.15 74.98 52.42 51.29 59.30
10 58.12 58.47 71.24 55.41 49.21 64.23
TABLE XI: NCF performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 89.31 0 96.18 0 83.23 0
2 85.95 16.64 93.12 12.37 81.93 27.43
3 82.51 19.62 91.36 15.59 77.59 30.73
4 79.98 21.39 88.93 19.25 73.90 33.19
5 72.45 24.83 85.13 23.82 70.20 37.32
6 69.12 28.27 81.09 29.13 67.14 43.79
7 67.78 35.45 77.42 36.66 62.85 47.41
8 64.90 39.10 75.33 41.55 57.22 53.33
9 59.72 44.85 71.98 46.32 50.59 58.20
10 58.12 50.97 69.24 52.41 48.31 63.23
TABLE XII: Siamese Networks performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 87.21 0 94.32 0 84.26 0
2 82.29 21.74 92.15 17.37 80.50 21.71
3 80.75 24.12 90.56 21.59 78.18 25.12
4 78.91 29.42 87.13 25.25 76.83 30.48
5 76.89 37.72 82.53 28.82 74.26 37.57
6 72.13 42.25 80.29 32.13 70.25 43.24
7 65.98 47.41 74.82 36.66 63.11 48.41
8 62.10 55.19 72.13 44.55 60.96 56.60
9 57.66 61.75 68.78 48.32 57.04 62.82
10 56.81 66.32 65.44 55.41 55.29 67.32
Figure 5: Accuracy and intra-list diversity visualization on top k recommendations
IV-F Interpretation of Results (RQ5)
The evaluation of the seven architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—reveals substantial performance variations based on dataset characteristics.
GNNs achieve the highest accuracy in the Retail E-commerce and Amazon datasets due to their ability to model complex graph structures. Transformers also perform strongly due to their capacity to extract global relationships through attention.
For the Netflix dataset, RNNs outperform others by effectively capturing sequential patterns in user rating histories, while GNNs remain robust but slightly less suited to temporal dynamics.
In terms of diversity, Siamese Networks offer the highest ILD on Retail Rocket due to their similarity-based optimization. CNNs lead in Amazon due to feature extraction variety, while GNNs yield the most diverse results for Netflix due to rich graph connectivity.
These results indicate that:
  • GNNs and Transformers are optimal for accuracy-focused recommendations.
  • Siamese Networks and CNNs provide strong diversity and can enhance hybrid recommenders.
  • The best model depends on the desired trade-off between accuracy and diversity.
IV-G Comparison with Other Works (RQ1, RQ2, RQ4)
Research on recommendation systems often relies on deep learning models such as matrix factorization techniques and conventional neural networks. Our approach, leveraging advanced models such as Graph Neural Networks (GNNs) and Transformers, brings a cutting-edge perspective by incorporating recent neural network developments. For instance, GNNs effectively model complex inter-item relationships, while Transformers utilize attention mechanisms, significantly enhancing traditional approaches.
While many conventional methods primarily focus on accuracy, our strategy emphasizes the diversity of recommendations. The Siamese Network, in particular, provides a unique contribution by emphasizing item-to-item similarities, broadening user engagement through more varied recommendations. This contrasts with typical systems that may overlook the importance of diversity on user satisfaction.
Our models also incorporate temporal dynamics, often underexplored in prior research. Recurrent Neural Networks (RNNs) excel in sequence prediction, allowing our systems to adapt recommendations to evolving user behavior. This adaptability is especially crucial for platforms like Netflix, where user preferences can change rapidly.
Across key metrics such as accuracy, recall, and F1-score, our models consistently demonstrate strong performance. The GNN model achieves superior results on the Retail Rocket and Amazon datasets, while the RNN model excels on the Netflix dataset. These outcomes surpass many existing solutions, providing more robust and reliable recommendations.
Furthermore, the effectiveness of our models in handling complex datasets, such as Amazon Products and Netflix Prize, demonstrates their robustness. These datasets contain a broad range of items and highly subjective user preferences, posing substantial challenges that our models successfully navigate.
Our comparative analysis highlights the potential for hybrid approaches that combine the strengths of various architectures to optimize both accuracy and diversity. Moreover, the results advocate for continued exploration of graph-based and attention-driven models in complex recommendation scenarios, with applications extending beyond e-commerce and media content.
Our models not only meet but advance current standards in recommendation system research. They provide a solid foundation for future efforts aimed at enhancing the scalability and adaptability of recommendation engines to diverse and dynamic market conditions.
IV-H Discussion on Limitations and Complications (RQ5)
Despite the advances achieved by our neural network models, several limitations and complications have been identified, highlighting areas for future improvement.
Computational Complexity: Advanced models like GNNs and Transformers require substantial computational resources due to their complex architectures. This can result in extended training times and higher operational costs, which may be impractical in environments with limited hardware or where rapid response is critical.
Dependency on Large, High-Quality Datasets: The models’ performance depends heavily on the availability of extensive and clean data. Sparse, noisy, or incomplete datasets can lead to reduced accuracy and biased recommendations. Additionally, reliance on large datasets raises concerns about user privacy and data security.
Balancing Accuracy and Diversity: Although our models perform well in complex environments, achieving a balance between recommendation accuracy and diversity remains challenging. Highly accurate recommendations may not always be sufficiently diverse, potentially limiting user engagement. Fine-tuning the models to recommend a broader range of items without compromising personalization is an ongoing challenge.
Adaptability to Real-World Dynamics: While our models incorporate temporal changes in user preferences, more dynamic real-time adaptation is needed to respond to sudden shifts in trends or user behavior. This is particularly relevant in fast-evolving domains such as fashion or entertainment, where user interests can change quickly.
While our models represent a significant advancement in recommendation systems, these limitations underscore the need for ongoing refinement. Future work should aim to improve computational efficiency, enhance data handling capabilities, and optimize the balance between diversity and relevance in real-world dynamic environments.
V Conclusion
Throughout this study, we have rigorously evaluated the performance of seven advanced neural network models across three diverse datasets—Retail E-commerce, Amazon Products, and Netflix Prize. Our findings demonstrate that these models, particularly Graph Neural Networks (GNNs) and Transformers, provide significant improvements over traditional recommendation systems by effectively handling complex data structures and emphasizing both accuracy and diversity in recommendations.
The GNN model excels in environments with intricate inter-item relationships, capturing the rich connectivity within user-item interactions. Transformers, with their efficient attention mechanisms, handle large datasets effectively, discerning nuanced relationships within sequences. Meanwhile, RNNs show superior performance in capturing temporal dynamics, making them especially suitable for platforms like Netflix, where user preferences evolve rapidly. This variation in model efficacy highlights the potential of tailored approaches that align with the specific characteristics and requirements of each dataset.
Despite these promising results, the study identifies several limitations, including computational demands, dependency on large, high-quality datasets, and challenges in balancing recommendation accuracy with diversity. These limitations underscore the need for ongoing research to optimize neural network architectures, improve computational efficiency, and enhance data processing strategies.
Looking ahead, the integration of hybrid models that combine the strengths of multiple architectures could lead to more robust, adaptable, and user-centric recommendation systems. Additionally, addressing ethical considerations related to data privacy and developing mechanisms capable of dynamically adjusting to real-world changes will be critical for the continued relevance of these systems.
This research not only advances our understanding of neural network applications in recommendation systems but also lays the groundwork for future innovations that can transform the way users interact with digital platforms. By continually refining these models, we can better meet the ever-changing preferences of users, ensuring that recommendation systems remain both relevant and impactful in the digital age.
Author Contributions
Abderaouf Bahi: Writing – original draft, Validation, Methodology, Investigation, Conceptualization, Visualization, Software. Ibtissem Gasmi: Writing – review and editing, Methodology, Supervision.
Acknowledgements
The authors acknowledge the Algerian Ministry of Higher Education and Scientific Research (MESRS).
Ethical Approval
Not Applicable.
Competing Interests
The authors declare no conflict of interest.
Funding
This research received no external funding.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.